Calculate the value of net national product (NNI) at market prices.
1
Expert's answer
2020-10-02T06:30:42-0400

use the information in the table below to answer Q.3.1 to Q.3.3

GDP at market price R397bn

net primary income payments for the rest of the world R37bn

indirect taxes R23bn

Subsidies R11bn

consumption of fixed capital R32bn

Answer

1 Calculate the value of gross national income [GNI] at market price

"GNP=GDP + net factor income from abroad"

"397+-37\n=360 bn"

2 Calculate the value of net national product (NNI) at market prices.

"NNP=GNP - depreciation\/ consumption of fixed capital"

"NNP =360-32=328 bn"